OpenStudy (anonymous):

A theater company charges $3.50 per ticket for people of all age groups. If c represents the total number of children who bought tickets and a represents the total number of adults who bought tickets, which expression best represents the total money earned by the company from the sale of tickets? 3.5c + a 3.5(c + a) c + 3.5a 7(c + a)

OpenStudy (anonymous):

@Jhannybean please help me!!!!!

OpenStudy (anonymous):

I just need the answer, please help me!

OpenStudy (jhannybean):

I can't give you the answer :( Sorry~

OpenStudy (jhannybean):

I could try working this out with you

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

No, B

OpenStudy (jhannybean):

C = children and A = adults

OpenStudy (anonymous):

Oh, so then it is B!!

OpenStudy (anonymous):

Right?

OpenStudy (jhannybean):

An expression for the total number of tickets the children bought is 3.50C (because each child pays $3.50) Total number of tickets bought by adults is 3.50A

OpenStudy (jhannybean):

So: 3.50C + 3.50A = 3.50(C+A)

OpenStudy (anonymous):

Thanks!!

OpenStudy (jhannybean):

no problem.

OpenStudy (anonymous):

Which unit rate is the lowest price per ounce? Choice A: 5 ounces of raisins for $1.49 Choice B: 12 ounces of raisins for $3.69 Choice A Choice B The unit rates are equal. The unit rates cannot be determined.

OpenStudy (anonymous):

Which one @Jhannybean?

OpenStudy (jhannybean):

I'm not too sure about this one.

OpenStudy (jhannybean):

Is this the entire question?

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

I guess you could just compare by dividing how how # of oz/price

OpenStudy (jhannybean):

To figure out the price for each.

OpenStudy (jhannybean):

choice A \(\dfrac{5~oz}{$1.49} = ~$3.35/1~oz\) choice B: \(\dfrac{12~oz}{$3.69} = $3.25/1~oz\)